Ane's brother Andrew leaves home for school at 8:00 am. He walks at 3.3 kph. At 8:20 am Jane discovers that Andrew has left his homework at home. She decides to follow him on her bicycle and give him his homework, but she wants to be back home in time for an online test at 9 am. What is the minimum constant speed at which she needs to ride? (Hint: At what time of day will they meet?)


Jane must ride at least ______ kph.


1
Expert's answer
2020-09-24T11:09:29-0400

"d=vt=4.2(t+\frac{20}{60})"


"2t=1h"


"0.5v=4.2(0.5+\frac{20}{60})= 7km/h"


Thus max speed= 7km/h.